# The Analogue of Aldous’ spectral gap conjecture for the generalized exclusion process

### v2

@@ 1.1. Background
The stationary distribution of this process is the function $\psi:S\rightarrow\mathbb{R}_{\geq 0}$ such that $\mathcal{L}\psi=\psi$. It is well known that when a continuous-time Markov process is reversible (i.e. $q_{x,y}\psi(x)=q_{y,x}\psi(y)$), all eigenvalues of $-\mathcal{L}$ are real and non-negative (c.f. Lemma 2.8). If such a continuous-time Markov process is irreducible, the eigenvalue $-\lambda_{i}$ for $i=0,\dots,|S|-1$ of $-\mathcal{L}$ satisfies
@@
In this paper, we evaluate the spectral gap of the following process. A complete symmetric weighted directed graph of size $n$ is a pair $X=(V,r)$ consisting of a set of vertices $V$ such that $|V|=n$ and a weight function on edges $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ such that $r(u,w)=r(w,u)$ and $r(u,u)=0$ for any $u,w\in V$. Fix $k\in\mathbb{Z}_{\geq 0}$, which we call the maximal occupancy, and fix $l\in\mathbb{Z}_{\geq 1}$, which we call the number of particles. We define $S_{GEP}=S_{GEP}(X,l,k)$ to be the set of assignments of $l$ indistinguishable particles to $n$ vertices, such that there are at most $k$ particles at each vertex. In other words,
$\begin{split}S_{GEP}:=\left\{\pi:V\rightarrow\{0,\dots,k\}\,\middle|\,\sum_{u\in V}\pi(u)=l\right\}.\end{split}$
For any state $\pi\in S_{GEP}$ and $u,v\in V$, if $\pi(u)>0$ and $\pi(v)<k$, we define $\pi^{uv}$ to be the state with one particle at vertex $u$ of state $\pi$ transitioning to vertex $v$ (cf. Figure1), that is
@@
For any $u,v\in V$, we assume the transition from a state $\pi\in S_{GEP}$ to $\pi^{uv}\in S_{GEP}$ occurs with rate $\mu(\pi,u,v)r(u,v)$, where $\mu:S_{GEP}\times V\times V\rightarrow\mathbb{R}_{\geq 0}$. Then, we define the generalized exclusion process as the Markov process with a state space $S_{GEP}$ and a generator
$\mathcal{L}^{GEP}f(\pi)=\sum_{u,v\in V}\mu(\pi,u,v)r(u,v)(f(\pi^{uv})-f(\pi)),$
where $f:S_{GEP}\rightarrow\mathbb{R}$ and $\pi\in S_{GEP}$. In this paper, we consider the case $\mu(\pi,u,v)=\pi(u)(k-\pi(v))$ for $u,v\in V$ and $\pi\in S_{GEP}$. We refer to a generalized exclusion process with such a transition rate as a normal generalized exclusion process
@@
The transition rate of a normal generalized exclusive process means that vertices with more particles are more likely to spit out particles and vertices with more empty spaces are more likely to accept particles. The normal generalized exclusion process is also an example of a generalized exclusion process that satisfies the gradient condition in a certain sense (c.f §3.2). The normal generalized exclusion process is a special case in which there is no “reservoir” of the process studied in [FGJS] and [JS].
@@
where $f:V\rightarrow\mathbb{R}$ and $u\in V$.
@@
$\lambda_{1}^{GEP}=k\lambda_{1}^{\mathrm{RW}}.$
@@
An interchange process on $V$ with rate $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ is a continuous-time Markov process with state space $S_{\mathrm{IP}}$ which gives the assignments of $n$ labeled particles to $V$ such that each vertex is occupied by exactly one particle. For any state $\eta\in S_{\mathrm{IP}}$, we let $\eta^{uv}$ denote a state in which the labels of vertices $u$ and $v$ of $\eta$ are swapped. We assume that a transition from $\eta\in S_{IP}$ to $\eta^{uv}$ occurs with rate $r(u,v)$. Then, the generator of the interchange process is defined by
@@ 1.2. Strategy
As studied by Cesi [Ces10], Aldous’ spectral gap conjecture was reformulated using Cayley graphs. Moreover, a generalization and analogue of Aldous’ spectral gap conjecture in the context of Cayley graphs has been studied by Parazanchevski and Puder [PP20], Li, Xia, and Zhou [LXZ23]. In §3, we describe the generalized exclusion process as the quotient of a Cayley graph associated with an interchange process. Finally, we apply the theorem to generalized exclusion processes and prove Theorem 1.3.
@@ 2.1. Complete weighted directed graph
If a complete weighted directed graph $\mathcal{S}$ is irreducible and reversible, by Perron-Frobenius theorem and Lemma 2.8, the eigenvalues $\lambda_{0},\lambda_{1},\dots,\lambda_{n}$ of $-\mathcal{L}^{\mathcal{S}}$ satisfies
$0=\lambda_{0}<\lambda_{1}<\cdots<\lambda_{n}.$
Then we can define the spectral gap as follows.
@@
To overcome the second part of this strategy, we apply Aldous’ spectral gap conjecture. In the next subsection, we will show how to reduce the proof of this equality to Aldous’ spectral gap conjecture.
@@ 2.2. Group action and quotient
The group $G$ acts on $\mathcal{S}$ if and only if we have $W(xg,y)=W(x,yg^{-1})$.
@@
Now, we show that a quotient of a weighted group induces a morphism of complete weighted graphs.
@@
$\begin{split}W_{G}(xG,yG)&=\sum_{g\in G}W(x,yg)\\ &=\sum_{g\in G/H}\sum_{h\in H}W(x,ygh)\\ &=\sum_{y^{\prime}H\in(yG)/H}W_{H}(xH,y^{\prime}H).\end{split}$
@@
However, as the last comment in §2.1 indicates, this does not give an effective answer for 1.1. We next prove an inequality obtained by a different approach and give a more effective answer to 1.1. This is a generalization of the discussion in [CLR10]*§4.1.1.
Theorem 2.16.
We assume that a finite group $G$ acts on $\mathcal{S}$ transitively (i.e. $xG=S$ for any $x\in S$). Let $H$ and $K$ be subgroups of $G$. We define $\mathcal{K}:=\{hKh^{-1}\mid h\in H\}$. We let $\check{K}$ denote $\bigcap_{K^{\prime}\in\mathcal{K}-\{K\}}K^{\prime}$. We assume the following two conditions;
  1. (1)
    The intersection $J=\cap_{K^{\prime}\in\mathcal{K}}K^{\prime}$ is a subgroup of $H$,
  2. (2)
    The subgroup of $G$ generated by $K$ and $\check{K}$ is equal to $G$.
Then we have
$\lambda_{1}^{\mathcal{S}/K}\geq\lambda_{1}^{\mathcal{S}/H}.$
To prove Theorem 2.16, we first prove the next lemma.
Lemma 2.17.
In the setting of Theorem 2.16, For any $h\in H$, the map
$\varphi_{h}:S/hKh^{-1}\rightarrow S/K;\quad xhKh^{-1}\mapsto xhK$
is well-defined. Moreover, this map $\varphi_{h}$ induces an isomorphism between $\mathcal{S}/hKh^{-1}$ and $\mathcal{S}/K$.
Proof.
For any $g\in K$, we have
$(xhgh^{-1})hKh^{-1}\mapsto xhgh^{-1}hK=xhgK.$
Therefore, the map $\varphi_{h}$ is well-defined and bijective. By Lemma 2.3, it suffices to show that $W_{K}(xhK,yK)=W_{hKh^{-1}}(xhKh^{-1},(yh^{-1})hKh^{-1})$ for any $x,y\in S$. This follows from the following calculation:
$\begin{split}W_{K}(xhK,yK)&=\sum_{g\in K}W(xh,yg)\\ &=\sum_{g\in K}W(x,yh^{-1}hgh^{-1})\\ &=W_{hKh^{-1}}(xhKh^{-1},(yh^{-1})hKh^{-1}).\end{split}$
Therefore, $\varphi_{h}$ gives an isomorphism of complete weighted directed graphs. ∎
Proof of Theorem 2.16.
We consider an eigenvector $f$ of $-\mathcal{L}^{\mathcal{S}/K}$ with eigenvalue $\lambda_{1}(\mathcal{S}/K)$. It suffices to show that $\lambda_{1}(\mathcal{S}/K)$ is also an eigenvalue of $-\mathcal{L}^{\mathcal{S}/H}$.
For any $K_{\alpha}\in\mathcal{K}$, we choose $h_{\alpha}\in H$ such that $h_{\alpha}Kh_{\alpha}^{-1}=K_{\alpha}$. If $K_{\alpha}=K$, we may assume that $h_{\alpha}$ is the identity element of $H$. For any $K_{\alpha}\in\mathcal{K}$, let $\varphi_{\alpha}:S/K_{\alpha}\rightarrow S/K$ be a map defined in Lemma 2.17. Then, this gives a function $f_{\alpha}:=f\circ\varphi_{\alpha}\in(\mathcal{S}/K_{\alpha})^{*}$ such that $f_{\alpha}(xK_{\alpha})=f(xh_{\alpha}K)$ for $xK_{\alpha}\in\mathcal{S}/K_{\alpha}$ is an eigenvector with eigenvalue $\lambda_{1}^{\mathcal{S}/K_{\alpha}}=\lambda_{1}^{\mathcal{S}/K}$. Then we define a function $f_{J}\in(\mathcal{S}/J)^{*}$ by
$f_{J}:=\sum_{K_{\alpha}\in\mathcal{K}}f_{\alpha}\circ\varphi_{J,\alpha}$
where $\varphi_{J,\alpha}$ is the natural projection $\mathcal{S}/J\rightarrow\mathcal{S}/K_{\alpha}$. If $K_{\alpha}=K$, we simply write $\varphi_{J}=\varphi_{J,\alpha}$.
We next prove that $f_{J}\neq 0$. If $f_{J}=0$, we have $f\circ\varphi_{J}=-\sum_{K_{\alpha}\in\mathcal{K}-\{K\}}f\circ\varphi_{J,\alpha}$. For any $x\in S$ and any $g\in K_{\alpha}$, we have
$f\circ\varphi_{J,\alpha}(xgJ)=f_{\alpha}(xgK_{\alpha})=f_{\alpha}(xK_{\alpha}).$
Therefore, for any $g\in\check{K}$, we have
$\begin{split}f(xgK)&=f\circ\varphi_{J}(xgJ)\\ &=-\sum_{K_{\alpha}\in\mathcal{K}-\{K\}}f_{\alpha}\circ\varphi_{J,\alpha}(xgJ)\\ &=-\sum_{K_{\alpha}\in\mathcal{K}-\{K\}}f_{\alpha}(xK_{\alpha})\\ &=f\circ\varphi_{J}(xJ)\\ &=f(xK).\end{split}$
Since a subgroup of $G$ generated by $K$ and $\check{K}$ is equal to $G$, for any $x\in S$ and $g\in G$, we have $f(xgK)=f(xK)$. Therefore, since $G$ acts on $\mathcal{S}$ transitively, we have $f$ is a constant function. However, this contradicts the assumption that $f$ is an eigenvector of $-\mathcal{L}^{\mathcal{S}/K}$ with eigenvalue $\lambda_{1}^{\mathcal{S}/K}>0$. Therefore, we have $f_{J}\neq 0$.
Since, for any $xJ\in\mathcal{S}/J$, we have
$\begin{split}-\mathcal{L}^{\mathcal{S}/J}f_{J}(xJ)&=-\mathcal{L}^{\mathcal{S}/J}\left(\sum_{K_{\alpha}\in\mathcal{K}}f_{\alpha}\circ\varphi_{J,\alpha}\right)(xJ)\\ &=-\sum_{K_{\alpha}\in\mathcal{K}}\mathcal{L}^{\mathcal{S}/J}(f_{\alpha}\circ\varphi_{J,\alpha})(xJ)\\ &=-\sum_{K_{\alpha}\in\mathcal{K}}\mathcal{L}^{\mathcal{S}/K_{\alpha}}f_{\alpha}(xK_{\alpha})\\ &=\sum_{K_{\alpha}\in\mathcal{K}}\lambda_{1}^{\mathcal{S}/K}f_{\alpha}(xK_{\alpha})\\ &=\lambda_{1}^{\mathcal{S}/K}f_{J}(xJ),\end{split}$
the map $f_{J}$ is an eigenvector of $-\mathcal{L}^{\mathcal{S}/J}$ with eigenvalue $\lambda_{1}^{\mathcal{S}/K}$. Let $\varphi:S/J\rightarrow S/H$ be the natural projection. Since, the functor $f_{H}:=f_{J}\circ\varphi$ is an eigenvector of $-\mathcal{L}^{\mathcal{S}/H}$ with eigenvalue $\lambda_{1}^{\mathcal{S}/K}$, we have $\lambda_{1}^{\mathcal{S}/K}\geq\lambda_{1}^{\mathcal{S}/H}$. ∎
@@ 3.1. Interchange Process
We note that the symmetric group $\mathfrak{S}_{n}$ naturally acts on $S_{\mathrm{IP}}$ from the left by the composition $\eta\cdot g:=g^{-1}\eta$ in $\mathfrak{S}_{n}$.
@@
The natural action of $\mathfrak{S}_{n}$ on $S_{\mathrm{IP}}$ induces an action of $\mathfrak{S}_{n}$ on $\mathcal{S}$.
@@ 3.2. Generalized exclusion process
Fix $k\in\mathbb{Z}_{\geq 1}$, which we call the maximal occupancy at $v$. We fix $l\in\mathbb{Z}_{\geq 1}$, which we call the number of particles, such that $l<N$. We define a set $S_{GEP}=S_{GEP}(X,l,k)$ as a set of assignments of $l$ indistinguishable particles to $n$ vertices such that each vertex has at most $k$ particles, that is
$\begin{split}S_{GEP}:=\left\{\pi:V\rightarrow\{0,1,\dots,k\}\,\middle|\,\sum_{u\in V}\pi(u)=l\right\}.\end{split}$
For any state $\pi\in S_{GEP}$ and $u,v\in V$, if $\pi(u)>0$ and $\pi(v)<k$, we define $\pi^{uv}$ to be the state with one particle at vertex $u$ of state $\pi$ transitioning to vertex $v$ (cf. Figure 3), that is
@@
For any $u,v\in V$, we assume a transition from a state $\pi\in S_{GEP}$ to $\pi^{uv}\in S_{GEP}$ occurs with rate $\mu(\pi,u,v)r(u,v)$ where $\mu:S_{GEP}\times V\times V\rightarrow\mathbb{R}_{\geq 0}$. Then, we define the generalized exclusion process as a continuous-time Markov process with state space $S_{GEP}$ and generator
$\mathcal{L}^{GEP}f(\pi)=\sum_{u,v\in V}\mu(\pi,u,v)r(u,v)(f(\pi^{uv})-f(\pi)),$
where $f:S_{GEP}\rightarrow\mathbb{R}$ and $\pi\in S_{GEP}$. In this paper, we consider the spectral gap of the generalized exclusion process with the special transition rate $\mu$ such that $\mu(\pi,u,v)=\pi(u)(k-\pi(v))$ for $u,v\in V$, $\pi\in S$. We refer to a generalized exclusion process with such a transition rate as a normal generalized exclusion process. We call a normal generalized exclusion process as a symmetric exclusion process if $k=1$. We call a symmetric exclusion process a random walk if $l=1$.
@@
The normal generalized exclusion process satisfies a gradient condition in the following sense (cf. [KL]*Remark 2.4). Let $V=\{u,v\}$, and a weight function $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ satisfy $r(u,v)=r(v,u)=1$. We define $f_{w}\in\mathrm{Map}(S_{GEP},\mathbb{R})$, $w\in\{u,v\}$ to be $f_{w}(\pi)=\pi(w)$. Then $f_{w}$, $w\in\{u,v\}$ is the conserved quantity of the generalized exclusion process. Moreover, we have
$\begin{split}-\mathcal{L}^{GEP}f_{u}(\pi)&=-\sum_{w_{1},w_{2}\in V}\mu(\pi,w_{1},w_{2})r(w_{1},w_{2})(\pi^{w_{1}w_{2}}(u)-\pi(u))\\ &=-\mu(\pi,u,v)r(u,v)(\pi^{uv}(u)-\pi(u))-\mu(\pi,v,u)r(v,u)(\pi^{vu}(u)-\pi(u))\\ &=-\pi(u)(k-\pi(v))((\pi(u)-1)-\pi(u))-\pi(v)(k-\pi(u))((\pi(u)+1)-\pi(u))\\ &=k(\pi(u)-\pi(v))\end{split}$
Thus, it can be said that $\mathrm{GEP}(X)$ satisfies the gradient condition.
@@
The main result of this paper is the proof of an analogue of Aldous’ spectral gap conjecture for the normal generalized exclusion process. More precisely, we prove that the spectral gap of the normal generalized exclusion process is equal to that of a random walk on a complete graph of the same order. To prove the main theorem, we will apply our framework and Aldous’ spectral gap conjecture. First, we will describe the normal generalized exclusion process as the quotient of the interchange process on the extended graph.
@@
Let $\widetilde{V}$ be a set $V\times\{1,\dots,k\}$, and let $\varphi_{V}:\widetilde{V}\rightarrow V$ be the natural projection. We define a map $\widetilde{r}:\widetilde{V}\rightarrow\mathbb{R}_{\geq 0}$ by
@@
$\begin{split}\widetilde{r}(\widetilde{u},\widetilde{v})&=\left\{\begin{aligned} {2}r(\varphi_{V}(\widetilde{u}),\varphi_{V}(\widetilde{v}))&&\mbox{if }\varphi_{V}(\widetilde{u})\neq\varphi_{V}(\widetilde{v}),\\ \sum_{w\in V-\varphi_{V}(\widetilde{u})}r(\varphi_{V}(\widetilde{u}),w)&&\mbox{if }\varphi_{V}(\widetilde{u})=\varphi_{V}(\widetilde{v}).\end{aligned}\right.\end{split}$
The complete weighted directed graph $\widetilde{X}=(\widetilde{V},\widetilde{r})$ is symmetric. (cf. Figure 4.)
Figure 4. The upper row of this picture shows the underlying graph and the generalized exclusion process on it, and the lower row shows the extended graph and the interchange process on this graph. In the picture of IP, we regard particles $1$ and $2$ as the particles of GEP.
@@
Let $\mathrm{IP}(\widetilde{X})=(\widetilde{S}_{\mathrm{IP}},\widetilde{W}_{\mathrm{IP}})$ be the complete weighted directed graph associated to the interchange process on $\widetilde{X}$. In the following, we will prove that the normal generalized exclusion process on $(X,r)$ is isomorphic to a quotient of $\mathrm{IP}(\widetilde{X})$.

### v3

@@ 1.1. Background
The stationary distribution of this process is the function $\psi:S\rightarrow\mathbb{R}_{\geq 0}$ such that $\mathcal{L}\psi=\psi$ and $\sum_{x\in S}\psi(x)=1$. It is well known that when a continuous-time Markov process is reversible (i.e. $q_{x,y}\psi(x)=q_{y,x}\psi(y)$), all eigenvalues of $-\mathcal{L}$ are real and non-negative (c.f. Lemma 2.10). If such a continuous-time Markov process is irreducible, that is every state can be reached from every other state (c.f [MC]*§1.2), the eigenvalue $-\lambda_{i}$ for $i=0,\dots,|S|-1$ of $-\mathcal{L}$ satisfies
@@
In this paper, we evaluate the spectral gap of the following process. A complete symmetric weighted directed graph of size $n$ is a pair $X=(V,r)$ consisting of a set of vertices $V$ such that $|V|=n$ and a weight function on edges $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ such that $r(u,w)=r(w,u)$ and $r(u,u)=0$ for any $u,w\in V$. Fix $k\in\mathbb{Z}_{\geq 0}$, which we call the maximal occupancy, and fix $l\in\mathbb{Z}_{\geq 1}$, which we call the number of particles. We define $S_{\mathrm{GEP}}=S_{\mathrm{GEP}}(X,l,k)$ to be the set of assignments of $l$ indistinguishable particles to $n$ vertices, such that there are at most $k$ particles at each vertex. In other words,
$\begin{split}S_{\mathrm{GEP}}:=\left\{\pi:V\rightarrow\{0,\dots,k\}\,\middle|\,\sum_{u\in V}\pi(u)=l\right\}.\end{split}$
For any state $\pi\in S_{\mathrm{GEP}}$ and $u,v\in V$ with $u\neq w$, if $\pi(u)>0$ and $\pi(v)<k$, we define $\pi^{uv}$ to be the state with one particle at vertex $u$ of state $\pi$ transitioning to vertex $v$ (cf. Figure1), that is
@@
For any $u,v\in V$, we assume the transition from a state $\pi\in S_{\mathrm{GEP}}$ to $\pi^{uv}\in S_{\mathrm{GEP}}$ occurs with rate $\mu(\pi,u,v)r(u,v)$, where $\mu:S_{\mathrm{GEP}}\times V\times V\rightarrow\mathbb{R}_{\geq 0}$. Then, we define the generalized exclusion process as the Markov process with a state space $S_{\mathrm{GEP}}$ and a generator
$\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}f(\pi)=\sum_{u,v\in V}\mu(\pi,u,v)r(u,v)(f(\pi^{uv})-f(\pi)),$
where $f:S_{\mathrm{GEP}}\rightarrow\mathbb{R}$ and $\pi\in S_{\mathrm{GEP}}$. In this paper, we consider the case $\mu(\pi,u,v)=\pi(u)(k-\pi(v))$ for $u,v\in V$ and $\pi\in S_{\mathrm{GEP}}$. We refer to a generalized exclusion process with such a transition rate as a normal generalized exclusion process
@@
The transition rate of a normal generalized exclusive process means that vertices with more particles are more likely to spit out particles and vertices with more empty spaces are more likely to accept particles. The normal generalized exclusion process is also an example of a generalized exclusion process that satisfies the gradient condition in a certain sense (c.f §3.2). The normal generalized exclusion process is a special case, in which there is no ”reservoir” as studied in the processes in [FGJS] and [JS], that is, there is no change in the total particle number.
@@
where $f:V\rightarrow\mathbb{R}$ and $u\in V$. We note that this is the special case of the generalized exclusion process with $k=l=1$.
@@
$\lambda_{1}^{\mathrm{GEP}_{k,l}(X)}=k\lambda_{1}^{\mathrm{RW}}.$
@@
An interchange process on $V$ with rate $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ is a continuous-time Markov process with state space $S_{\mathrm{IP}}$ which gives the assignments of $n$ labeled particles to $V$ such that each vertex is occupied by exactly one particle. For any state $\eta\in S_{\mathrm{IP}}$, we let $\eta^{uv}$ denote a state in which the labels of vertices $u$ and $v$ of $\eta$ are swapped. We assume that a transition from $\eta\in S_{\mathrm{IP}}$ to $\eta^{uv}$ occurs with rate $r(u,v)$. Then, the generator of the interchange process is defined by
@@
Remark 1.7.
Our main result Theorem 1.3 for the special case $k=1$ and $l=n$ follows directly from Theorem 1.6.
@@ 1.2. Strategy
As studied by Cesi [Ces10], Aldous’ spectral gap conjecture was reformulated using Cayley graphs. Moreover, a generalization and analogue of Aldous’ spectral gap conjecture in the context of Cayley graphs has been studied by Parazanchevski and Puder [PP20], Li, Xia, and Zhou [LXZ23]. In §2, we study the spectral gap of complete weighted directed graphs and their quotients. We will restate the results of classical continuous-time Markov processes in our setting. In §3, we describe the generalized exclusion process as the quotient of a Cayley graph associated with an interchange process. Finally, we apply the theorem to generalized exclusion processes and prove Theorem 1.3.
@@ 2.1. Complete weighted directed graph
Example 2.3.
Let $\mathcal{S}_{1}=(S_{1},W_{1})$ be a weighted graph on $S_{1}=\{1,2,3\}$ with the weight
$W_{1}(x,y)=\begin{cases}1&x\neq y\\ 0&x=y.\end{cases}$
We let $\mathcal{S}_{2}=(S_{2},W_{2})$ be a weighted graph on $S_{2}:=\{(x,i)\mid x\in S_{1},i=1,2\}$ with the weight
$W_{2}((x,i),(y,j))=\begin{cases}1&x\neq y\,\mathrm{and}\,i\neq j\\ 0&\mathrm{otherwise}.\end{cases}$
Then the map $\varphi_{12}:S_{2}\rightarrow S_{1}$ such that $\varphi_{12}((x,i))=x$ is an morphism $\varphi_{12}:\mathcal{S}_{2}\rightarrow\mathcal{S}_{1}$.
Example 2.4.
Let $\mathcal{S}_{3}=(S_{3},W_{3})$ be a weighted graph on $S_{3}=\{1,2\}$ with the weight
$W_{3}(x,y)=\begin{cases}2&x\neq y\\ 0&x=y.\end{cases}$
We let $\mathcal{S}_{4}=(S_{4},W_{4})$ be a weighted graph on $S_{4}:=\{(x,i)\mid x\in S_{3},i=1,2\}$ with the weight
$W_{4}((x,i),(y,j))=\begin{cases}1&x\neq y\\ 1&x=y.\end{cases}$
Then the map $\varphi_{34}:S_{4}\rightarrow S_{3}$ such that $\varphi_{34}((x,i))=x$ is an morphism $\varphi_{34}:\mathcal{S}_{4}\rightarrow\mathcal{S}_{3}$.
Figure 2. The left column represents the morphism of Example 2.3, and the right column represents the morphism of Example 2.4.
@@
If a complete weighted directed graph $\mathcal{S}=(S,W)$ is irreducible and reversible, by Perron-Frobenius theorem and Lemma 2.10, the eigenvalues $\lambda_{0},\lambda_{1},\dots,\lambda_{n}$ of $-\mathcal{L}^{\mathcal{S}}$ satisfies
$0=\lambda_{0}<\lambda_{1}\leq\cdots\leq\lambda_{n-1},$
where $n=|S|$. Then we can define the spectral gap as follows.
@@
In this paper, we apply Aldous’s spectral gap conjecture to overcome the obstacles in the second part of this strategy. In the next subsection, we will show how to reduce the proof of this equality to Aldous’ spectral gap conjecture.
@@ 2.2. Group action and quotient
We assume that $G$ acts on a weighted graph $(S,W)$, that is for any $g\in G$ and $x,y\in S$, we have $W(xg,yg)=W(x,y)$. Then, it induces an action on $\mathcal{S}$ if and only if we have $W(xg,y)=W(x,yg^{-1})$ for any $x,y\in S$ and $g\in G$.
Proof.
We define the bijection $\varphi_{g}:S\mapsto S$ by $\varphi_{g}(x):=xg$. Then, by Lemma 2.5, the bijection $\varphi_{g}$ induces the automorphism on $\mathcal{S}$ if and only if
$W(xg,y)=W(\varphi_{g}(x),y)=\sum_{y^{\prime}\in\varphi_{g}^{-1}(y)}W(x,y^{\prime})=W(x,yg^{-1}).$
for any $x,y\in S$ and $g\in G$. ∎
@@
Now, we show that a quotient of a weighted graph induces a morphism of complete weighted graphs.
@@
$\begin{split}W_{G}(xG,yG)&=\sum_{g\in G}W(x,yg)\\ &=\sum_{gH\in G/H}\sum_{h\in H}W(x,ygh)\\ &=\sum_{y^{\prime}H\in(yG)/H}W_{H}(xH,y^{\prime}H).\end{split}$
@@ 3.1. Interchange Process
We note that the symmetric group $\mathfrak{S}_{n}$ naturally acts on $S_{\mathrm{IP}}$ from the right by the composition $\eta\cdot g:=g^{-1}\eta$ in $\mathfrak{S}_{n}$.
@@
The natural action of $\mathfrak{S}_{n}$ on $S_{\mathrm{IP}}$ induces an action of $\mathfrak{S}_{n}$ on $\mathrm{IP}(X)$.
@@ 3.2. Generalized exclusion process
Fix $k\in\mathbb{Z}_{\geq 1}$, which we call the maximal occupancy at $v$. We fix $l\in\mathbb{Z}_{\geq 1}$, which we call the number of particles, such that $l<N$. We define a set $S_{\mathrm{GEP}}=S_{\mathrm{GEP}}(X,l,k)$ as a set of assignments of $l$ indistinguishable particles to $n$ vertices such that each vertex has at most $k$ particles, that is
$\begin{split}S_{\mathrm{GEP}}:=\left\{\pi:V\rightarrow\{0,1,\dots,k\}\,\middle|\,\sum_{u\in V}\pi(u)=l\right\}.\end{split}$
For any state $\pi\in S_{\mathrm{GEP}}$ and $u,v\in V$, if $\pi(u)>0$ and $\pi(v)<k$, we define $\pi^{uv}$ to be the state with one particle at vertex $u$ of state $\pi$ transitioning to vertex $v$ (cf. Figure 4), that is
@@
For any $u,v\in V$, we assume a transition from a state $\pi\in S_{\mathrm{GEP}}$ to $\pi^{uv}\in S_{\mathrm{GEP}}$ occurs with rate $\mu(\pi,u,v)r(u,v)$ where $\mu:S_{\mathrm{GEP}}\times V\times V\rightarrow\mathbb{R}_{\geq 0}$. Then, we define the generalized exclusion process as a continuous-time Markov process with state space $S_{\mathrm{GEP}}$ and generator
$\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}f(\pi)=\sum_{u,v\in V}\mu(\pi,u,v)r(u,v)(f(\pi^{uv})-f(\pi)),$
where $f:S_{\mathrm{GEP}}\rightarrow\mathbb{R}$ and $\pi\in S_{\mathrm{GEP}}$. In this paper, we consider the spectral gap of the generalized exclusion process with the special transition rate $\mu$ such that $\mu(\pi,u,v)=\pi(u)(k-\pi(v))$ for $u,v\in V$, $\pi\in S$. We refer to a generalized exclusion process with such a transition rate as a normal generalized exclusion process. We call a normal generalized exclusion process as a symmetric exclusion process if $k=1$. We call a symmetric exclusion process a random walk if $l=1$.
@@
The normal generalized exclusion process satisfies a gradient condition in the following sense (cf. [KL]*Remark 2.4). Let $V=\{u,v\}$, and a weight function $r:V\times V\rightarrow\mathbb{R}_{\geq 0}$ satisfy $r(u,v)=r(v,u)=1$. We define $f_{w}\in\mathrm{Map}(S_{\mathrm{GEP}},\mathbb{R})$, $w\in\{u,v\}$ to be $f_{w}(\pi)=\pi(w)$. Then $f_{w}$, $w\in\{u,v\}$ is the conserved quantity of the generalized exclusion process. Moreover, we have
$\begin{split}-\mathcal{L}^{\mathrm{GEP}_{k,l}(X)}f_{u}(\pi)&=-\sum_{w_{1},w_{2}\in V}\mu(\pi,w_{1},w_{2})r(w_{1},w_{2})(\pi^{w_{1}w_{2}}(u)-\pi(u))\\ &=-\mu(\pi,u,v)r(u,v)(\pi^{uv}(u)-\pi(u))-\mu(\pi,v,u)r(v,u)(\pi^{vu}(u)-\pi(u))\\ &=-\pi(u)(k-\pi(v))((\pi(u)-1)-\pi(u))-\pi(v)(k-\pi(u))((\pi(u)+1)-\pi(u))\\ &=k(\pi(u)-\pi(v))\end{split}$
Thus, it can be said that $\mathrm{GEP}_{k,l}(X)$ satisfies the gradient condition.
@@
The main result of this paper is the proof of an analogue of Aldous’ spectral gap conjecture for the normal generalized exclusion process. More precisely, we explicitly presents the spectral gap of the normal generalized exclusion process by $k$-times of that of a random walk on a complete graph of the same order. To prove the main theorem, we will apply our framework and Aldous’ spectral gap conjecture. First, we will describe the normal generalized exclusion process as the quotient of the interchange process on an extended graph.
@@
Let $\widetilde{V}$ be the set $V\times\{1,\dots,k\}$, and let $\varphi_{V}:\widetilde{V}\rightarrow V$ be the natural projection. We define a map $\widetilde{r}:\widetilde{V}\rightarrow\mathbb{R}_{\geq 0}$ by
@@
$\begin{split}\widetilde{r}(\widetilde{u},\widetilde{v})&=\left\{\begin{aligned} r(\varphi_{V}(\widetilde{u}),\varphi_{V}(\widetilde{v}))&&\mbox{if }\varphi_{V}(\widetilde{u})\neq\varphi_{V}(\widetilde{v}),\\ \sum_{w\in V-\varphi_{V}(\widetilde{u})}r(\varphi_{V}(\widetilde{u}),w)&&\mbox{if }\varphi_{V}(\widetilde{u})=\varphi_{V}(\widetilde{v}).\end{aligned}\right.\end{split}$
The complete weighted directed graph $\widetilde{X}=(\widetilde{V},\widetilde{r})$ is symmetric.
@@
Let $\mathrm{IP}(\widetilde{X})=(\widetilde{S}_{\mathrm{IP}},\widetilde{W}_{\mathrm{IP}})$ be the complete weighted directed graph associated with the interchange process on $\widetilde{X}$. In the following, we will prove that the normal generalized exclusion process on $(X,r)$ is isomorphic to a quotient of $\mathrm{IP}(\widetilde{X})$.